# New affine invariant ensemble samplers and their dimensional scaling

### v2

@@ 4.3. Hamiltonian walk move sampler
will also preserve the joint distribution. A similar antisymmetric preconditioning has been explored in [35] for underdamped Langevin dynamics.
@@
  • We fix particles in $S^{(1)}$, and update particles in $S^{(0)}$: for each $21\leq i\leq N/$ 2, we sample $\mathbf{p}_{i}\sim\mathcal{N}(0,I_{N/2\times N/2})$, and run $n$ steps of leapfrog approximation of the dynamics
@@
    We propose $(\tilde{\mathbf{x}}_{i},\tilde{\mathbf{p}}_{i})=\mathsf{P}\mathsf{L}_{h,B_{S^{(1)}}}^{n}(\mathbf{x}_{i},\mathbf{p}_{i})$, and we accept this proposal for each $21\leq i\leq N/$ 2 with probability
@@
(4.10) $(\mathbf{x}_{1},...,\mathbf{x}_{N})\overset{\phi}{\to}(\mathbf{y}_{1},...,\mathbf{y}_{N})=(A\mathbf{x}_{1}+\mathbf{b},...,A\mathbf{x}_{N}+\mathbf{b})$
@@
We note that the centered ensembles here play a role similar to the Cholesky factor of the empirical covariance matrix. While it is possible to directly use the Cholesky factor for preconditioning, which would result in a momentum vector with the same dimension as the state, using centered ensembles avoids the computational cost of factorization. Similar use of centered ensembles to avoid Cholesky has been considered in [22] for covariance preconditioned Langevin dynamics.
The algorithm may be interpreted as a gradient-enhanced walk move. In one form of the walk move [26], the proposal is $\mathbf{x}_{i}+B_{S^{(1)}}\mathbf{z}$ with $\mathbf{z}\sim\mathcal{N}(0,I_{N/2\times N/2})$. In our algorithm, the derivative of $\mathbf{x}_{i}$ is similarly a linear combination of the columns of $B_{S^{(1)}}$. When the integration time is small, the leading order term follows the same direction as $B_{S^{(1)}}\mathbf{z}$. With longer integration times, our approach adaptively determines the move strength using Hamiltonian-type dynamics that incorporate derivative information of the potential. For this reason, we term the method the Hamiltonian walk move sampler. In general, similar to the walk move sampler, we can choose an arbitrary subset $S$ from the complementary ensemble to form the centered ensemble; in the above Hamiltonian walk move algorithm, $S$ is chosen to contain all particles in the complementary ensemble. Choosing a smaller $S$ could reduce the arithmetic cost.
@@ 4.4. Hamiltonian side move sampler
  • We fix particles in $S^{(1)}$, and update particles in $S^{(0)}$: for each $21\leq i\leq N/$ 2, we sample $p_{i}\sim\mathcal{N}(0,1)$ and two particles $\mathbf{x}_{j}$ and $\mathbf{x}_{k}$ from $S^{(1)}$. We run $n$ steps of leapfrog approximation of the dynamics
@@ 5.1. Evaluation criterion
We investigate the efficiency of affine invariant ensemble samplers through numerical experiments. Our main evaluation criterion is the autocorrelation time at the stationary phase, following [26]. These samplers generate sequences $(\mathbf{x}_{1}(m),...,\mathbf{x}_{N}(m))$ for $1\leq m\leq M$, where $M$ represents the length of the ensemble chain. We use these ensembles to estimate the observable
@@
where $\tau_{s}=\sum_{m=-\infty}^{+\infty}\frac{C_{s}(m)}{C_{s}(0)}$ and $C_{s}(m)=\lim_{m^{\prime}\to\infty}\mathrm{Cov}[f(\mathbf{x}(m^{\prime})),f(\mathbf{x}(m+m^{\prime}))]$. As noted by [26], a natural criterion to compare performance is the values of $\tau_{e}$ and $\tau_{s}$. The estimation procedure of the autocorrelation time follows [26, Section 5].
@@ 5.2. Synthetic: Gaussian
Figure 3. Rescaled autocorrelation functions for sampling anisotropic Gaussian targets with condition number $\kappa=1000$. Left: stretch move; right: side move. Rescaled lag $=$ original lag///dim×$\times 4$.
@@
Figure 4. Autocorrelation time (thinning by $10$) versus dimension for anisotropic Gaussian targets with condition number $\kappa=1000$. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog steps.
@@ 5.4. Invariant distribution to stochastic PDEs
In this illustration, we generate samples from an infinite-dimensional probability measure defined over continuous functions on the unit interval $[0,1]$; see [26]. The measure is formally
@@
subject to natural boundary conditions at both endpoints $x=0$ and $x=1$. Here, $\eta$ denotes space-time white noise. Realizations from this distribution typically exhibit rough, approximately constant profiles near either $1$ or $-1$; thus this is a bimodal distribution. We discretize the derivatives using finite difference with equidistributed points. Denote the total number of points by $d$, which leads to a $d$ dimensional distribution.
We adopt the same setup as before. The autocorrelation function is computed for the path integral observable $f(u)=\int_{0}^{1}u(x){\rm d}x$ discretized using composite trapezoid rules. In Figure 5, we show the scaling of the autocorrelation time. We observe similar $O(d)$ scaling for stretch and side moves. The HMC scales worse than in the Gaussian example, since the condition number of the SPDE example deteriorates as dimension grows and we use a fixed number of leapfrog steps. For the affine invariant Hamiltonian walk move, we observe a nearly constant autocorrelation time as before, which demonstrates its superior performance in sampling such ill-conditioned distributions. In particular, for $d=64$, Hamiltonian walk move with $n=2$ leapfrog steps leads to 100×100$\times$100 × reduction in autocorrelation time compared to HMC with $n=10$ leapfrog steps [^2] In this example with $d=64$, the acceptance rate for HMC is $0.12$, while for the Hamiltonian walk move it is $0.70$..
@@
We note that in this test example, the target distribution is absolutely continuous with respect to a Gaussian measure; accordingly, function space MCMC [16] is also expected to achieve dimension-robust convergence. Combining affine invariant samplers with function space MCMC may further enhance performance [17, 19].
Figure 5. Autocorrelation time (thinning by $10$) versus dimension for the bimodal SPDE targets. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog steps.
@@ 6. Discussions and Conclusions
Algorithmically, adapting the step size and, in the case of affine invariant HMC, the number of leapfrog steps could further improve performance, according to the wide success of the No-U-Turn Sampler (NUTS) [29, 8]; in particular, there has been some recent developments on tuning-free multichain adaptive HMC methods [30, 28, 52, 51]. In our experiments, we observe that the Hamiltonian side move leads to smaller autocorrelation time compared to the standard side move. In each iteration, they move in the same direction, suggesting potential for further adaptation of the move length of the derivative-free side move. A recent relevant development is the No-Underrun Sampler (NURS) [5], which focuses on adapting move lengths along random line directions without using derivatives.
Theoretically, establishing the ergodicity of these affine invariant ensemble samplers remains an important challenge. To the best of our knowledge, the only work in this direction is [22] for covariance-preconditioned Langevin dynamics. Still, their analysis applies to continuous dynamics rather than discrete algorithms. Moreover, in addition to the autocorrelation time at the stationary phase, understanding the mixing time or convergence during the burn-in period is also of interest. In this direction, existing results [21, 9] primarily focus on Gaussian target distributions in continuous-time covariance-preconditioned Langevin dynamics. We also note that affine invariant algorithms projected onto the Gaussian family (in the sense of variational inference) has been shown to achieve exponential improvement in condition number dependence when the target is a one-dimensional log-concave distribution [12, Theorem 5.7].
Finally, beyond affine invariance, there have been developments leveraging diffeomorphism invariant dynamics for sampling, particularly through the Fisher-Rao gradient flow [12]. The numerical approximation of diffeomorphism-invariant dynamics is subtle, and several algorithms can be interpreted as implementations of the Fisher-Rao gradient flow; these include methods based on birth-death processes [38, 39], Kalman methodology and variational inference [31, 13, 10], and ensemble chains [36]. Typically, at most affine invariance is preserved in the numerical implementation and the samplers are often approximate and have bias. It is an interesting direction to integrate these advances to develop exact or approximate samplers with broader invariance properties.
Acknowledgments YC is grateful to Jonathan Goodman for many intellectual discussions and shared insights on MCMC that inspired this work. YC also thanks Andrew Stuart and Jonathon Weare for valuable feedback on an earlier version of this article.
@@ Appendix A Pseudocode for All the Affine Invariant Samplers
Algorithm 1 The parallel stretch move at step $m$ (adapted from [20])
0:  Ensemble $S^{(s)}(m)$ for $s\in\{0,1\}$, number of walkers $N$, unnormalized probability density $\pi(\cdot)$, parameter $a$
@@
0:  Ensemble $S^{(s)}(m)$ for $s\in\{0,1\}$, number of walkers $N$, unnormalized probability density $\pi(\cdot)$, parameter $\sigma$
@@
0:  Ensemble $S^{(s)}(m)$ for $s\in\{0,1\}$, number of walkers $N$, unnormalized probability density $\pi(\cdot)$, leapfrog stepsize $h$ and steps $n$
@@
Algorithm 4 The parallel Hamiltonian side walk move at step $m$
0:  Ensemble $S^{(s)}(m)$ for $s\in\{0,1\}$, number of walkers $N$, unnormalized probability density $\pi(\cdot)$, leapfrog stepsize $h$ and steps $n$
@@ B.2. Scaling of stretch move
$g(z)\propto\begin{cases}\frac{1}{\sqrt{z}}&\text{if }z\in\left[\frac{1}{a},a\right]\\ 0&\text{otherwise}\,,\end{cases}$
@@
converges in distribution to $(\beta U,z_{1},z_{2},\delta_{1},\delta_{0})$ (for convergence in distribution in the product space, see [4, Section 2]). Here $U\sim\mathrm{Unif}[-1,1],z_{1}\sim\mathcal{N}(0,3),z_{2}\sim\mathcal{N}(0,1)$ are independent, and $\delta_{x}$ is a Diracs point pass at $x$. The convergence is obtained using the law of large numbers and the central limit theorem as well as the fact that $Z$ and $\mathbf{x}_{i},\mathbf{x}_{j}$ are independent. We also used the fact that convergence almost surely implies convergence in distribution.
@@ C.1. A simple formula for the acceptance probability
This formula is similar to the one derived in [1] for standard Hamiltonian dynamics. We adapt the strategy to derive the formula for the above preconditioned Hamiltonian dynamics. Here, the additional $B$ requires us to discuss the rank of the matrix $B^{T}B$.

### v3

@@ 4.3. Hamiltonian walk move sampler
will also preserve the joint distribution. A similar antisymmetric preconditioning has been explored in [leimkuhler2018ensemble] for underdamped Langevin dynamics.
@@
  • We fix particles in $S^{(1)}$, and update particles in $S^{(0)}$: for each $1\leq i\leq N/2$, we sample $\mathbf{p}_{i}\sim\mathcal{N}(0,I_{N/2\times N/2})$, and run $n$ steps of leapfrog approximation of the dynamics
@@
    We propose $(\tilde{\mathbf{x}}_{i},\tilde{\mathbf{p}}_{i})=\mathsf{P}\mathsf{L}_{h,B_{S^{(1)}}}^{n}(\mathbf{x}_{i},\mathbf{p}_{i})$, and we accept this proposal for each $1\leq i\leq N/2$ with probability
@@
(4.10) $(\mathbf{x}_{1},.,\mathbf{x}_{N})\overset{\phi}{\to}(\mathbf{y}_{1},.,\mathbf{y}_{N})=(A\mathbf{x}_{1}+\mathbf{b},.,A\mathbf{x}_{N}+\mathbf{b})$
@@
We note that the centered ensembles here play a role similar to the Cholesky factor of the empirical covariance matrix. While it is possible to directly use the Cholesky factor for preconditioning, which would result in a momentum vector with the same dimension as the state, using centered ensembles avoids the computational cost of factorization. Similar use of centered ensembles to avoid Cholesky has been considered in [garbuno2020affine] for covariance preconditioned Langevin dynamics.
The algorithm may be interpreted as a gradient-enhanced walk move. In one form of the walk move [goodman2010ensemble], the proposal is $\mathbf{x}_{i}+B_{S^{(1)}}\mathbf{z}$ with $\mathbf{z}\sim\mathcal{N}(0,I_{N/2\times N/2})$. In our algorithm, the derivative of $\mathbf{x}_{i}$ is similarly a linear combination of the columns of $B_{S^{(1)}}$. When the integration time is small, the leading order term follows the same direction as $B_{S^{(1)}}\mathbf{z}$. With longer integration times, our approach adaptively determines the move strength using Hamiltonian-type dynamics that incorporate derivative information of the potential. For this reason, we term the method the Hamiltonian walk move sampler. In general, similar to the walk move sampler, we can choose an arbitrary subset $S$ from the complementary ensemble to form the centered ensemble; in the above algorithm, $S$ is chosen to contain all particles in the complementary ensemble. Choosing a smaller $S$ could reduce the arithmetic cost.
@@ 4.4. Hamiltonian side move sampler
  • We fix particles in $S^{(1)}$, and update particles in $S^{(0)}$: for each $1\leq i\leq N/2$, we sample $p_{i}\sim\mathcal{N}(0,1)$ and two particles $\mathbf{x}_{j}$ and $\mathbf{x}_{k}$ from $S^{(1)}$. We run $n$ steps of leapfrog approximation of the dynamics
@@ 5.1. Evaluation criterion
We investigate the efficiency of affine invariant ensemble samplers through numerical experiments. Our main evaluation criterion is the autocorrelation time at the stationary phase, following [goodman2010ensemble]. These samplers generate sequences $(\mathbf{x}_{1}(m),...,\mathbf{x}_{N}(m))$ for $1\leq m\leq M$, where $M$ represents the length of the ensemble chain. We use these ensembles to estimate the observable
@@
where $\tau_{s}=\sum_{m=-\infty}^{+\infty}\frac{C_{s}(m)}{C_{s}(0)}$ and $C_{s}(m)=\lim_{m^{\prime}\to\infty}\mathrm{Cov}[f(\mathbf{x}(m^{\prime})),f(\mathbf{x}(m+m^{\prime}))]$. As noted by [goodman2010ensemble], a natural criterion to compare performance is the values of $\tau_{e}$ and $\tau_{s}$. The estimation procedure of the autocorrelation time follows [goodman2010ensemble, Section 5].
@@ 5.2. Synthetic: Gaussian
[…]
@@
Figure 4. Autocorrelation time (thinning by $10$) versus dimension for anisotropic Gaussian targets with condition number $\kappa=1000$. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$, and the number of leapfrog steps is $n=10$ for HMC while $n=2$ for affine invariant HMC.
@@ 5.4. Invariant distribution to stochastic PDEs
In this illustration, we generate samples from an infinite-dimensional probability measure defined over continuous functions on the unit interval $[0,1]$; see [goodman2010ensemble]. The measure is formally
@@
subject to natural boundary conditions at both endpoints $x=0$ and $x=1$. Here, $\eta$ denotes space-time white noise. Realizations from this distribution typically exhibit rough, approximately constant profiles near either $1$ or 1-1; thus this is a bimodal distribution. We discretize the derivatives using finite difference with equidistributed points. Denote the total number of points by $d$, which leads to a $d$ dimensional distribution.
We adopt the same setup as before. The autocorrelation function is computed for the path integral observable $f(u)=\int_{0}^{1}u(x){\rm d}x$ discretized using composite trapezoid rules. In Figure 5, we show the scaling of the autocorrelation time. We observe similar $O(d)$ scaling for stretch and side moves. The HMC scales worse than in the Gaussian example, since the condition number of the SPDE example deteriorates as dimension grows and we use a fixed number of leapfrog steps. For the affine invariant Hamiltonian walk move, we observe a nearly constant autocorrelation time as before, which demonstrates its superior performance in sampling such ill-conditioned distributions. In particular, for $d=64$, Hamiltonian walk move with $n=2$ leapfrog steps leads to $100\times$ reduction in autocorrelation time compared to HMC with $n=10$ leapfrog steps [^2] In this example with $d=64$, the acceptance rate for HMC is $0.12$, while for the Hamiltonian walk move it is $0.70$..
@@
We note that in this test example, the target distribution is absolutely continuous with respect to a Gaussian measure; accordingly, function space MCMC [cotter2013mcmc] is also expected to achieve dimension-robust convergence. Combining affine invariant samplers at the coarse scale with function space MCMC at the fine scale may further enhance performance [coullon2021ensemble, dunlop2022gradient].
Figure 5. Autocorrelation time (thinning by $10$) versus dimension for the bimodal SPDE targets. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$, and the number of leapfrog steps is $n=10$ for HMC while $n=2$ for affine invariant HMC.
@@ 6. Discussions and Conclusions
Algorithmically, adapting the step size and, in the case of affine invariant HMC, the number of leapfrog steps could further improve performance, according to the wide success of the No-U-Turn Sampler (NUTS) [hoffman2014no, carpenter2017stan]; in particular, there has been some recent developments on tuning-free multichain adaptive HMC methods [hoffman2022tuning, hoffman2021adaptive, sountsov2021focusing, sountsov2024running]. In our experiments, we observe that the Hamiltonian side move leads to smaller autocorrelation time compared to the standard side move. In each iteration, they move in the same direction, suggesting potential for further adaptation of the move length of the derivative-free side move. A recent relevant development is the No-Underrun Sampler (NURS) [bou2025no], which focuses on adapting move lengths along random line directions without using derivatives.
Theoretically, establishing the ergodicity of these affine invariant ensemble samplers remains an important challenge. To the best of our knowledge, the only work in this direction is [garbuno2020affine] for covariance-preconditioned Langevin dynamics. Still, their analysis applies to continuous dynamics rather than discrete algorithms. Moreover, in addition to the autocorrelation time at the stationary phase, understanding the mixing time or convergence during the burn-in period is also of interest. In this direction, existing results [garbuno2020interacting, carrillo2021wasserstein] primarily focus on Gaussian target distributions in continuous-time covariance-preconditioned Langevin dynamics. We also note that affine invariant algorithms projected onto the Gaussian family (in the sense of variational inference) has been shown to achieve exponential improvement in condition number dependence when the target is a one-dimensional log-concave distribution [chen2023sampling, Theorem 5.7].
Finally, beyond affine invariance, there have been developments leveraging diffeomorphism invariant dynamics for sampling, particularly through the Fisher-Rao gradient flow [chen2023sampling]. The numerical approximation of diffeomorphism-invariant dynamics is subtle, and several algorithms can be interpreted as implementations of the Fisher-Rao gradient flow; these include methods based on birth-death processes [lu2019accelerating, lu2023birth], Kalman methodology and variational inference [huang2022efficient, chen2024efficient, che2025stable], and ensemble chains [lindsey2022ensemble]. Typically, at most affine invariance is preserved in the numerical implementation and the samplers are often approximate and have bias. It is an interesting direction to integrate these advances to develop exact or approximate samplers with broader invariance properties.
Acknowledgments YC is grateful to Jonathan Goodman for intellectual discussions on MCMC that inspired this work. YC also thanks Andrew Stuart and Jonathon Weare for helpful communications.
@@ Appendix A Pseudocode for All the Affine Invariant Samplers
Algorithm 1 The parallel stretch move at step $m$ (adapted from [foreman2013emcee])
0: Ensemble $S^{(s)}(m)$ for $s\in\{0,1\}$, number of walkers NN, unnormalized probability density $\pi(\cdot)$, parameter $a$
@@
0: Ensemble $S^{(s)}(m)$ for $s\in\{0,1\}$, number of walkers NN, unnormalized probability density $\pi(\cdot)$, parameter $\sigma$
@@
0: Ensemble $S^{(s)}(m)$ for $s\in\{0,1\}$, number of walkers NN, unnormalized probability density $\pi(\cdot)$, leapfrog stepsize $h$ and steps $n$
@@
Algorithm 4 The parallel Hamiltonian side move at step $m$
0: Ensemble $S^{(s)}(m)$ for $s\in\{0,1\}$, number of walkers NN, unnormalized probability density $\pi(\cdot)$, leapfrog stepsize $h$ and steps $n$
@@ B.2. Scaling of stretch move
[…]
@@
converges in distribution to $(\beta U,z_{1},z_{2},\delta_{1},\delta_{0})$ (for convergence in distribution in the product space, see [billingsley2013convergence, Section 2]). Here $U\sim\mathrm{Unif}[-1,1],z_{1}\sim\mathcal{N}(0,3),z_{2}\sim\mathcal{N}(0,1)$ are independent, and $\delta_{x}$ is a Diracs point pass at $x$. The convergence is obtained using the law of large numbers and the central limit theorem as well as the fact that $Z$ and $\mathbf{x}_{i},\mathbf{x}_{j}$ are independent. We also used the fact that convergence almost surely implies convergence in distribution.
@@ C.1. A simple formula for the acceptance probability
This formula is similar to the one derived in [apers2024hamiltonian] for standard Hamiltonian dynamics. We adapt the strategy to derive the formula for the above preconditioned Hamiltonian dynamics. Here, the additional $B$ requires us to discuss the rank of the matrix $B^{T}B$.